Justin wants to have at least a 75 test aveerage in Algebra 1 this six weeks. He will have 3 test grades. He recieved a 72 and a 76 on the first two tests. What must he make on the third test to receive at least a 75 average?
Answer by stanbon(75887)

Let x be the score on the third test.
The average he wants must be [72+76+x]/3>=75
72+76+x>=225
148+x>=225
x>=77
Cheers,
Stan H.